# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 1. Introduction
on $\mathbb{T}^{d}\times[0,T]$ for some $T>0$ and any spatial dimension $d\geq 2$, equipped with divergence-free initial data at $t=0$. The real parameter $\nu>0$ represents viscosity, which we typically normalize to $1$ without loss of generality. We say that a solution $(u,p)$ of (1.1) is classical if it is smooth on $\mathbb{T}^{d}\times[0,T]$ in space and time. The pioneering work of Leray [Leray1934] established local-in-time existence of classical solutions from any smooth initial data, as well as global existence of weak solutions to (1.1) in any dimension $d\geq 2$, for any finite-energy initial data.
@@
When $d=2$, finite-time blow-up was ruled out in the whole space by Leray [Leray1934] and in bounded domains by Ladyzhenskaya [Ladyzhenskaya1967]. For $d\geq 3$, the question is open. Demonstration of a finite-time blow-up for $d=3$ would, by uniqueness, contradict global existence, resolving the Millennium Prize Problem on global regularity for the 3D Navier–Stokes equations in the negative direction.
@@
An alternative perspective on instantaneous blow-up is as a finite-time blow-up that proceeds backwards in time, in the sense that norms of the solution become large as $t\to T_{*}+$. One might expect such a blow-up to occur more easily than the conventional forward-in-time blow-up because it would be facilitated by the Laplacian. We emphasize that this is not the case. Indeed, many of the critical regularity criteria (for instance, Ladyzhenskaya–Prodi–Serrin and Onsager-type ones) that are well-known to prevent finite-time blow-up also prevent instantaneous blow-up. In fact, the time asymmetry at the heart of the construction is from not the Laplacian, but the nonlinearity. Specifically, the key observation is that nonlocal interactions in frequency space are most robust from high to low frequencies forwards in time, or equivalently, from low to high frequencies backwards in time.
If a smooth solution of Navier–Stokes undergoes an instantaneous blow-up at $t=T_{*}$ as defined above, failure of uniqueness and even weak-strong uniqueness of the initial value problem directly follow. Indeed, because the solution is smooth at the blow-up time, one may apply standard local well-posedness theory with initial data $u(T_{*})$ to obtain a classical solution $U$ that agrees with $u$ on $0,T_{*}]$, but remains uniformly bounded in a neighborhood of the blow-up time. By (1.3), $u$ and $U$ must be distinct after $T_{*}$. In fact, in the examples constructed here, something much stronger is true: the classical solution $U$ is the limit of an infinite family of solutions that exhibit instantaneous blow-up at $t=T_{*}$ (see Figure 2). The family consists of quasi-periodic solutions on the whole space that exhibit translation symmetry-breaking in the sense that the solutions spontaneously lose periodicity. A discrete subset of the family consists of genuinely periodic solutions with the same period as the background classical solution. The solutions we exhibit blow up in the Type I sense (see Subsection 1.3 for definitions and discussion) and lie arbitrarily close to known criteria for uniqueness and weak-strong uniqueness of Navier–Stokes weak solutions.
@@ 1.1. Main results
We emphasize that, by definition, classical solutions are smooth in space and time. In particular, combining Properties 12, the solution $u(t)\in C^{\infty}(\mathbb{T}^{d})$ for every $t\in[0,T]$, including at the blow-up time. In fact, such solutions are analytic in space and time away from $t=T_{*}$; see [MR245989, MR870865]. See Figure 1 for a schematic depiction of the regularity of the solution.
@@
Figure 1. Blow-up from the right.
@@
With $T^{u_{0}}\in(0,\infty]$ the maximum time of existence for the classical solution with initial data $u_{0}$, the time $T$ in Theorem 1.1 can be taken as any $T<T^{u_{0}}$. One can apply the theorem repeatedly to construct $u$ on the full lifespan $0,T^{u_{0}})$, although one loses uniform estimates in the case $T^{u_{0}}<\infty$.
@@
The type of blow-up demonstrated in Theorem 1.1 does not appear to have been identified before for the Navier–Stokes equations or other fluid models. As the closest available precedent, we point to the instantaneous loss of regularity phenomenon appearing in work of Córdoba, Martínez-Zoroa, and Ożański for the 2D Euler equations [MR4776418] and the fractionally dissipative SQG equations [MR4800679]. They construct solutions with initially non-smooth data, say in $H^{1+\epsilon}$, which immediately drop down to lower regularity $H^{s}$ for some $s<1+\epsilon$. See also Section 1.6 for another comparison.
@@
$u^{(\sigma)}\big|_{[0,T_{*}]}\equiv U\big|_{[0,T_{*}]},\qquad u^{(\sigma)}\big|_{(T_{*},T]}\not\equiv U\big|_{(T_{*},T]}\qquad\forall\sigma\in(0,1].$
@@
$u^{(\sigma)}(t)\in C^{\infty}(\mathbb{R}^{d})\qquad\forall t\in[0,T]$
@@
We make the following technical observation: because the $u^{(\sigma)}$ are in $L_{t,x}^{\infty}$ on $\mathbb{R}^{d}\times[T^{\prime},T]$ for any $T^{\prime}>T_{*}$, there is no difficulty formulating the notions of weak and classical solutions. Indeed, in the framework of bounded mild solutions, the standard local theory carries over from $\mathbb{T}^{d}$ or decaying solutions on $\mathbb{R}^{d}$; see for instance [MR2545826].
@@
In this section, we expound on the significance of the blow-up constructed in Theorem 1.1 and its relevance to the broader question of well-posedness of the Navier–Stokes equations. Both types of smooth blow-up described above require a cascade of energy or some higher regularity norm across infinitely many frequency scales. They are differentiated by the direction of the cascade: whether the energy is transferred from low- to high-frequency modes or vice versa characterizes the blow-up type as finite-time or instantaneous, respectively. While finite-time blow-up has received far more attention, both phenomena are relevant for understanding the global behavior of Navier–Stokes solutions, and can conceivably be exhibited by Leray–Hopf solutions.
@@
The scenario (B) in Figure 3, in which both finite-time and instantaneous blow-up occur at $T_{*}$, is consistent with the Leray–Hopf energy inequality as long as the energy anomalously lost in the first blow-up exceeds the energy gained in the second. Even more strongly [^1] It is presently unknown whether every Leray–Hopf solution is a restricted Leray–Hopf solution., it is consistent with the known behavior of a restricted Leray–Hopf solution, i.e., the weak limit of a sequence of solutions of Galerkin-truncated Navier–Stokes, as it undergoes a potential blow-up. This highlights the fact that an instantaneous blow-up can occur in the Leray–Hopf class only after the appearance of a finite-time blow-up. In contrast, for general weak solutions, an instantaneous blow-up can occur from an arbitrary classical solution, as demonstrated in Theorem 1.1. While there are numerous regularity conditions precluding a finite-time blow-up, the analogous problem for instantaneous blow-up does not appear to have been considered. In Section LABEL:sec:regularity, we report some results to the effect that classical regularity criteria for finite-time blow-up can be extended to the instantaneous scenario.
@@
if $\limsup_{t\to T_{*}+}\|u(t)\|_{L^{\infty}}=\infty$, then necessarily the initial data is singular. An instantaneous blow-up arising from smooth initial data can occur only through a nonlinear mechanism that transfers energy across infinitely-many frequency modes.
@@
Finally, the instantaneous blow-up and non-uniqueness constructed in this paper involves a purely temporal loss of smoothness, where the solutions may evolve from any smooth initial data, while staying analytic in space for all $t>0$ and in time for all $t\neq T_{*}$. It therefore represents a fundamentally different manifestation of non-uniqueness—one that occurs entirely within the class of smooth, divergence-free vector fields (see Section 1.4 for further discussion.)
@@ 1.3. Criticality and Type I blow-up
for any $d<p<\infty$. Another important family of critical spaces is given by the Ladyzhenskaya–Prodi–Serrin scale, namely $L^{p}_{t}L^{q}_{x}$ where $q>3$ and $\frac{2}{p}+\frac{d}{q}=1$, with scale-invariance enforced by the latter condition.
@@
$\|u(t)\|_{L^{\infty}}\leq\frac{C}{\sqrt{T_{*}-t}},\quad t<T_{*}$
@@
$\|u(t)\|_{L^{\infty}}\leq\frac{C}{\sqrt{t-T_{*}}},\quad t>T_{*}$
@@
$\liminf_{t\to T_{*}+}\,\sqrt{t-T_{*}}\|u(t)\|_{L^{\infty}}\geq c$ (1.8)
for some $c>0$ by Property 2. Let us also point out that while there is not a single agreed upon definition of Type I blow-up, the definition given above appears to be the traditional one, first appearing in the context of the Ricci flow [MR1375255] and later for other types parabolic and dispersive equations ([MR2077706, MR2545826, MR3272053], etc.). In the recent fluids literature, it is common to refer to any bound obeyed by self-similar solutions as Type I, particularly those of the form $u\in L_{t}^{\infty}X$ where $X$ is in $L^{d,\infty}$ or any of the spaces in which it is embedded in (1.7). Under this definition, the instantaneous blow-up constructed here is still of Type I, in that the solution is in $L_{t}^{\infty}\dot{W}_{x}^{-1,\infty}$ or $L^{\infty}_{t}BMO^{-1}_{x}$.
For comparison, the existence of finite-time Type I blow-up for the Navier–Stokes equations with $d\geq 3$ is a major open question, although it has been ruled out in special cases including backward self-similar solutions [MR1397564, MR1643650], axisymmetric solutions [MR2545826, MR2429247, MR2512859, MR4628936], and when the critical norm is strengthened to, say, $L^{d}$ [Escauriaza].
@@ 1.4. Previous non-uniqueness constructions
For finite-energy initial data, Leray’s construction yields the existence of weak solutions satisfying the energy inequality (Leray–Hopf solutions), and possessing additional desirable properties, such as Leray’s structure theorem and the weak-strong uniqueness. In the classical setting, uniqueness holds under certain regularity or integrability conditions. Most notably, strong solutions, that is, solutions satisfying the Ladyzhenskaya–Prodi–Serrin condition
@@ 1.4.1. Convex integration
The Euler equations, by which we mean (1.1) with $\nu=0$, offer insight into the uniqueness problem. Following the pioneering works of Scheffer [Scheffer1993] and Shnirelman [Shnirelman1997, Shnirelman2000], the convex integration framework brought to mathematical fluid mechanics by De Lellis and Székelyhidi [DeLellisSzekelyhidi2009, DeLellisSzekelyhidi2013] provided a new powerful tool for constructing weak solutions of PDEs with various pathological properties motivated by physical theories of turbulence. This program led to Isett’s construction of non-conservative $C_{x,t}^{1/3-}$ solutions for the Euler equations in 3D [Isett2018], later extended to 2D by Giri and Radu in [MR4809443]. Together with the proof of energy conservation above the $\frac{1}{3}$-Hölder threshold [MR1298949], these works are regarded to have resolved Onsager’s conjecture on anomalous energy dissipation.
These developments exposed a profound breakdown of uniqueness and the classical notion of determinism in fluid mechanics. Still, the weak solutions of the Euler equations arising from convex integration are not known to be vanishing-viscosity limits of Navier–Stokes solutions in any reasonable regularity class, raising the question of which “wild” phenomena can be transferred to the viscous setting. Indeed, the Laplacian term in (1.1) provides a smoothing mechanism that might have the potential to preclude the wild oscillatory behavior observed in the Euler case. A breakthrough in this direction came in the work of Buckmaster and Vicol [BuckmasterVicol2019], who introduced a viscous adaptation of convex integration and constructed the first examples of non-unique weak solutions of (1.1). Their solutions lie in $H^{\epsilon}(\mathbb{T}^{3})$ for some $\epsilon>0$ and exhibit strong non-uniqueness and non-conservation of the $L^{2}$ norm. Subsequent constructions of wild solutions with improvements to the regularity were given, for instance, in [MR4422213, 2009.06596, MR4610908]. These results demonstrate that viscosity does not, in general, restore uniqueness. The recent constructions in [CZZ2025, CoiculescuPalasek2025] push this point further, as the Laplacian term is essential to the emergence of non-uniqueness, and the constructions fail in the inviscid case.

### v2

@@ 1. Introduction
on $\mathbb{T}^{d}\times[0,T]$ for some $T>0$ and any spatial dimension $d\geq 2$, equipped with divergence-free initial data at $t=0$. The real parameter $\nu>0$ represents viscosity, which we typically normalize to $1$ without loss of generality. We say that a solution $(u,p)$ of (1.1) is classical if it is smooth on $\mathbb{T}^{d}\times[0,T]$ in space and time. The pioneering work of Leray [36] established local-in-time existence of classical solutions from any smooth initial data, as well as global existence of weak solutions to (1.1) in any dimension $d\geq 2$, for any finite-energy initial data.
1.1. An inquiry into blow-up phenomena
@@
When $d=2$, finite-time blow-up was ruled out in the whole space by Leray [36] and in bounded domains by Ladyzhenskaya [33]. For $d\geq 3$, the question is open. Demonstration of a finite-time blow-up for $d=3$ would, by uniqueness, contradict global existence, resolving the Millennium Prize Problem on global regularity for the 3D Navier–Stokes equations in the negative direction.
@@
An alternative perspective on instantaneous blow-up is as a finite-time blow-up that proceeds backwards in time, in the sense that norms of the solution become large as $t\to T_{*}+$. One might expect such a blow-up to occur more easily than the conventional forward-in-time blow-up because it would be facilitated by the Laplacian. We emphasize that this is not the case. Indeed, many of the critical regularity criteria (for instance, Ladyzhenskaya–Prodi–Serrin and Onsager-type ones) that are well-known to prevent finite-time blow-up also prevent instantaneous blow-up. In fact, the time asymmetry at the heart of the construction is not from the Laplacian, but the nonlinearity. Specifically, the key observation is that nonlocal interactions in frequency space are most robust from high to low frequencies forwards in time, or equivalently, from low to high frequencies backwards in time.
If a smooth solution of Navier–Stokes undergoes an instantaneous blow-up at $t=T_{*}$ as defined above, failure of uniqueness and even weak–strong uniqueness of the initial value problem directly follow. Indeed, because the solution is smooth at the blow-up time, one may apply standard local well-posedness theory with initial data $u(T_{*})$ to obtain a classical solution $U$ that agrees with $u$ on $0,T_{*}]$, but remains uniformly bounded in a neighborhood of the blow-up time. By (1.3), $u$ and $U$ must be distinct after $T_{*}$. In fact, in the examples constructed here, something much stronger is true: the classical solution $U$ is the limit of an infinite family of solutions that exhibit instantaneous blow-up at $t=T_{*}$ (see Figure 2). The family consists of quasi-periodic solutions on the whole space that exhibit translation symmetry-breaking in the sense that the solutions spontaneously lose periodicity. A discrete subset of the family consists of genuinely periodic solutions with the same period as the background classical solution. The solutions we exhibit blow up in the Type I sense (see Subsection 2.2 for definitions and discussion) and lie arbitrarily close to known criteria for uniqueness and weak–strong uniqueness of Navier–Stokes weak solutions.
@@ 1.2. Main results
We emphasize that, by definition, classical solutions are smooth in space and time. In particular, combining Properties 12, the solution $u(t)\in C^{\infty}(\mathbb{T}^{d})$ for every $t\in[0,T]$, including at the blow-up time. In fact, such solutions are analytic in space and time away from $t=T_{*}$; see [29, 30]. See Figure 1 for a schematic depiction of the regularity of the solution.
@@
Figure 1. Instantaneous blow-up at $t=T_{*}$.
@@
The solution $u$ lies in neither $L_{t}^{\infty}L_{x}^{2}$ nor $L_{t}^{2}H_{x}^{1}$ near the blow-up time. We emphasize that an instantaneous blow-up of the energy is not a limitation of the construction, but in fact a necessary condition for instantaneous Type I blow-up of the $L^{\infty}$ norm; see Theorem 2.4.
@@
With $T^{u_{0}}\in(0,\infty]$, the maximum time of existence for the classical solution with initial data $u_{0}$, the time $T$ in Theorem 1.1 can be taken as any $T<T^{u_{0}}$. One can apply the theorem repeatedly to construct $u$ on the full lifespan $0,T^{u_{0}})$, although one loses uniform estimates in the case $T^{u_{0}}<\infty$.
Remark 1.6.
The type of blow-up demonstrated in Theorem 1.1 does not appear to have been identified before for the Navier–Stokes equations or other fluid models. As the closest available precedent, we point to the instantaneous loss of regularity phenomenon appearing in work of Córdoba, Martínez-Zoroa, and Ożański for the 2D Euler equations [14] and the fractionally dissipative SQG equations [15]. They construct solutions with initially non-smooth data, say in $H^{1+\epsilon}$, which immediately drop down to lower regularity $H^{s}$ for some $s<1+\epsilon$. Also, there are examples of self-similar non-unique solutions to semilinear equations [23]. See also Section 2.5 for another comparison.
@@
$u^{(\sigma)}\big|_{[0,T_{*}]}\equiv U\big|_{[0,T_{*}]},\qquad u^{(\sigma)}\big|_{(T_{*},T]}\not\equiv U\big|_{(T_{*},T]},\qquad\forall\sigma\in(0,1].$
@@
$u^{(\sigma)}(t)\in C^{\infty}(\mathbb{R}^{d}),\qquad\forall t\in[0,T],$
@@
We make the following technical observation: because the $u^{(\sigma)}$ are in $L_{t,x}^{\infty}$ on $\mathbb{R}^{d}\times[T^{\prime},T]$ for any $T^{\prime}>T_{*}$, there is no difficulty formulating the notions of weak and classical solutions. Indeed, in the framework of bounded mild solutions, the standard local theory carries over from $\mathbb{T}^{d}$ or decaying solutions on $\mathbb{R}^{d}$; see for instance [31].
Structure of the paper. We further explain the significance of the main results and highlight the main ideas of the construction in Section 2. Section 3 introduces the notations to be used in the paper. Sections 4, 5 and 6 contain the technical construction and estimates. Section 7 is devoted to the proof of the main results. Auxiliary lemmas are collected in the Appendix.
Acknowledgments. M.D. is partially supported by the National Science Foundation grant DMS-2308208 and the Simons Foundation. S.P. acknowledges support from the National Science Foundation under Grant No. DMS-2424441.
2. Significance of the results and main ideas
In this section, we expound on the significance of the blow-up constructed in Theorem 1.1 and non-uniqueness established in Theorem 1.7, and their relevance to the broader question of well-posedness of the Navier–Stokes equations.
@@ 2. Significance of the results and main ideas
Both types of smooth blow-up as described in Section 1.1 require a cascade of energy or some higher regularity norm across infinitely many frequency scales. They are differentiated by the direction of the cascade: whether the energy is transferred from low- to high-frequency modes or vice versa characterizes the blow-up type as finite-time or instantaneous, respectively. While finite-time blow-up has received far more attention, both phenomena are relevant for understanding the global behavior of Navier–Stokes solutions, and can conceivably be exhibited by Leray–Hopf solutions.
@@
The scenario (B) in Figure 3, in which both finite-time and instantaneous blow-up occur at $T_{*}$, is consistent with the Leray–Hopf energy inequality as long as the energy anomalously lost in the first blow-up exceeds the energy gained in the second. Even more strongly [^1] It is presently unknown whether every Leray–Hopf solution is a restricted Leray–Hopf solution., it is consistent with the known behavior of a restricted Leray–Hopf solution, i.e., the weak limit of a sequence of solutions of Galerkin-truncated Navier–Stokes, as it undergoes a potential blow-up. This highlights the fact that an instantaneous blow-up can occur in the Leray–Hopf class only after the appearance of a finite-time blow-up. In contrast, for general weak solutions, an instantaneous blow-up can occur from an arbitrary classical solution, as demonstrated in Theorem 1.1. While there are numerous regularity conditions precluding a finite-time blow-up, the analogous problem for instantaneous blow-up does not appear to have been considered.
@@
if $\limsup_{t\to T_{*}+}\|u(t)\|_{L^{\infty}}=\infty$, then necessarily the initial data is singular. An instantaneous blow-up arising from smooth initial data can occur only through a nonlinear mechanism that transfers energy across infinitely many frequency modes.
@@
Finally, the instantaneous blow-up and non-uniqueness constructed in this paper involve a purely temporal loss of smoothness, where the solutions may evolve from any smooth initial data, while staying analytic in space for all $t>0$ and in time for all $t\neq T_{*}$. It therefore represents a fundamentally different manifestation of non-uniqueness—one that occurs entirely within the class of smooth, divergence-free vector fields (see Section 2.3 for further discussion.)
@@ 2.2. Criticality and Type I blow-up
for any $d<p<\infty$. Another important family of critical spaces is given by the Ladyzhenskaya–Prodi–Serrin scale, namely $L^{p}_{t}L^{q}_{x}$ where $q>d$ and $\frac{2}{p}+\frac{d}{q}=1$, with scale-invariance enforced by the latter condition.
@@
$\|u(t)\|_{L^{\infty}}\leq\frac{C}{\sqrt{T_{*}-t}},\quad t<T_{*},$
@@
$\|u(t)\|_{L^{\infty}}\leq\frac{C}{\sqrt{t-T_{*}}},\quad t>T_{*},$
@@
$\liminf_{t\to T_{*}+}\,\sqrt{t-T_{*}}\|u(t)\|_{L^{\infty}}\geq c,$ (2.2)
for some $c>0$ by Property 2. Let us also point out that while there is not a single agreed upon definition of Type I blow-up, the definition given above appears to be the traditional one, first appearing in the context of the Ricci flow [22] and later for other types of parabolic and dispersive equations ([38, 31, 16], etc.). In the recent fluids literature, it is common to refer to any bound obeyed by self-similar solutions as Type I, particularly those of the form $u\in L_{t}^{\infty}X$ where $X$ is in $L^{d,\infty}$ or any of the spaces in which it is embedded in (2.1). Under this definition, the instantaneous blow-up constructed here is still of Type I, in that the solution is in $L_{t}^{\infty}\dot{W}_{x}^{-1,\infty}$ or $L^{\infty}_{t}BMO^{-1}_{x}$.
For comparison, the existence of finite-time Type I blow-up for the Navier–Stokes equations with $d\geq 3$ is a major open question, although it has been ruled out in special cases including backward self-similar solutions [42, 49], axisymmetric solutions [31, 6, 5, 43], and when the critical norm is strengthened to, say, $L^{d}$ [17].
@@ 2.3. Previous non-uniqueness constructions
For finite-energy initial data, Leray’s construction yields the existence of weak solutions satisfying the energy inequality (Leray–Hopf solutions), and possessing additional desirable properties, such as Leray’s structure theorem and the weak–strong uniqueness. In the classical setting, uniqueness holds under certain regularity or integrability conditions. Most notably, strong solutions, that is, solutions satisfying the Ladyzhenskaya–Prodi–Serrin condition
@@ 2.3.1. Convex integration
The Euler equations, by which we mean (1.1) with $\nu=0$, offer insight into the uniqueness problem. Following the pioneering works of Scheffer [46] and Shnirelman [47, 48], the convex integration framework brought to mathematical fluid mechanics by De Lellis and Székelyhidi [34, 35] provided a new powerful tool for constructing weak solutions of PDEs with various pathological properties motivated by physical theories of turbulence. This program led to Isett’s construction of non-conservative $C_{x,t}^{1/3-}$ solutions for the Euler equations in 3D [26], later extended to 2D by Giri and Radu in [20]. Together with the proof of energy conservation above the $\frac{1}{3}$-Hölder threshold [13], these works are regarded to have resolved Onsager’s conjecture on anomalous energy dissipation.
These developments exposed a profound breakdown of uniqueness and the classical notion of determinism in fluid mechanics. Still, the weak solutions of the Euler equations arising from convex integration are not known to be vanishing-viscosity limits of Navier–Stokes solutions in any reasonable regularity class, raising the question of which “wild” phenomena can be transferred to the viscous setting. Indeed, the Laplacian term in (1.1) provides a smoothing mechanism that might have the potential to preclude the wild oscillatory behavior observed in the Euler case. A breakthrough in this direction came in the work of Buckmaster and Vicol [3], who introduced a viscous adaptation of convex integration and constructed the first examples of non-unique weak solutions of (1.1). Their solutions lie in $H^{\epsilon}(\mathbb{T}^{3})$ for some $\epsilon>0$ and exhibit strong non-uniqueness and non-conservation of the $L^{2}$ norm. Subsequent constructions of wild solutions with improvements to the regularity were given, for instance, in [4, 9, 10]. These results demonstrate that viscosity does not, in general, restore uniqueness. The recent constructions in [11, 12] push this point further, as the Laplacian term is essential to the emergence of non-uniqueness, and the constructions fail in the inviscid case.